80.0 cm3 of hydrogen is collected during a chemistry experiment at 20.0°C and 101 kPa.


What is its volume at IUPAC stp

1
Expert's answer
2022-01-02T18:24:44-0500

The volume is


"V=\dfrac{TP_iV_i}{PT_1}=\dfrac{(273+0)101·80}{101.325(273+20)}=74\text{ cm}^3."